# Solar Photovoltaic Assessment with Large Language Model

### v1

@@ 1 Introduction
The global transition toward renewable energy has significantly influenced the operation and planning of power systems, particularly in active distribution networks (ADNs) and microgrids [1]. As a leading source of clean energy, solar photovoltaic (PV) technology has become increasingly important for decentralizing power generation and enhancing grid resilience [2]. To realize these benefits, accurate detection and mapping of solar PV installations in satellite imagery are critical for efficient energy management, demand-response optimization, and the seamless integration of distributed energy resources (DERs) [3]. However, the widespread adoption of solar panels has led to an exponential increase in the amount of imagery data, necessitating highly scalable and adaptable approaches for PV identification.
Early approaches to solar panel detection from imagery relied on traditional machine learning models, such as Logistic Regression [4], Random Forest [5], and Support Vector Machines based methods [6], often employing handcrafted features like texture and color.
@@
Although these methods were relatively efficient and required fewer training samples, they lacked the capacity to generalize to diverse spatial patterns. More recent research has turned towards deep learning [7, 8], particularly Convolutional Neural Network (CNN)-based models [9], which have demonstrated superior performance in extracting vision features from imagery data. Notable efforts such as Deep Solar [10] showcased large-scale solar panel mapping using Inception-v3, while U-Net [11] enabled global solar mapping. However, these methods often require substantial annotated datasets [12, 13], intensive computational resources, and frequent retraining to adapt to changing conditions in different geographical regions and environmental contexts.
@@
These limitations highlight a pressing need for flexible, transparent, and resource-efficient solutions. Large language models (LLMs) have emerged as a promising paradigm for detection tasks with limited labels [14], owing to their extensive pre-training on diverse data sources. Although originally developed for natural language processing (NLP), advanced LLMs such as OpenAI’s GPT series [15], Google’s PaLM [16], and Meta’s LLaMA [17] have extended their capabilities to multimodal tasks, including image analysis. LLMs can capture rich contextual information [18, 19, 20] and handle heterogeneous inputs and varying contexts, making them particularly suitable for the dynamic environments typical of ADNs and microgrids [21].
@@
Although LLMs show great promise for solar panel detection, they face significant challenges. Firstly, LLMs often struggle with multi-step logical processes essential for accurately identifying and assessing solar panels [22]. This difficulty arises because solar panel detection typically requires sequential reasoning, such as distinguishing panels from similar structures, understanding spatial relationships, and integrating contextual information from satellite imagery. Additionally, LLMs can exhibit inconsistencies in output formatting [23]. This inconsistency can lead to increased manual intervention, reducing the overall efficiency and scalability of solar energy management workflows.
@@
Another major challenge is the misclassification of visually similar objects. For example, shadows, parking lots, or other flat surfaces may be incorrectly identified as solar panels due to their comparable appearance in satellite images. This issue is exacerbated in environments with high visual complexity or varying lighting conditions, which can obscure or distort the features LLMs rely on for accurate classification. Furthermore, LLMs often display low accuracy in complex tasks such as spatial localization and quantification of solar panels. Accurate spatial localization requires precise mapping of panel locations, while quantification involves estimating the number and capacity of panels, both of which demand high precision and reliability.
@@
To address the challenges associated with LLMs in solar panel detection, we propose prompting and fine-tuning strategies. First, solar panel detection involves analyzing complex rooftop images to accurately identify and assess solar installations. To manage this complexity, we implement task decomposition within our PVAL framework. We begin by defining the overall objective in the prompt, guiding the model to understand the context and goals of the analysis. The task is then broken down into specific steps: first, the model performs a holistic image analysis to detect potential solar panels; second, it pinpoints the exact locations of these panels within predefined regions; and third, it estimates the quantity of panels based on their arrangement and visibility. This structured approach ensures that each sub-task is handled systematically, reducing the likelihood of errors and improving the overall accuracy of detection. By decomposing the task, we enable the LLM to focus on manageable components, thereby enhancing its ability to process complex visual data effectively.
@@
Second, consistent and interpretable outputs are crucial for integrating detection results into existing solar energy management systems. To achieve this, we employ output standardization by defining a JSON-based response format. This standardized structure includes key fields such as “solar panels present”, “location”, and “quantity”, each with predefined possible values to ensure uniformity. For example, the “location” field uses a detailed classification schema that divides the image into specific regions like “top-left” or “center” as illustrated in our visual schema (Figure 1). This standardization addresses the problem of inconsistent output formats, facilitating seamless data integration into geographic information systems (GIS) and energy management platforms. By ensuring that all outputs adhere to a uniform format, we eliminate the need for additional preprocessing, thereby streamlining workflows and enhancing the scalability of our detection system.
@@
Third, accurate classification of solar panels, especially when distinguishing them from visually similar objects like shadows or parking lots, is a significant challenge. To address this issue, we utilize few-shot prompting by incorporating a limited number of highly relevant examples within the prompt. These examples include both positive cases (where solar panels are present) and negative cases (where they are absent), providing the LLM with clear references on how to handle different scenarios. For instance, one example might specify the detection of solar panels in the “top-left” quadrant with a quantity range of “0 to 1”, while another example indicates no panels detected by marking “location” and “quantity” as “NA”. By using five-shot prompting, we enhance the model’s ability to generalize from these examples, improving its classification accuracy without the need for extensive labeled datasets. This approach mitigates misclassification issues, enabling the LLM to better distinguish between similar-looking objects and thereby increasing the reliability of solar panel detection.
@@
Last, another challenge for LLMs in solar panel detection is handling complex spatial localization and quantification tasks. Accurate spatial localization requires precise mapping of panel locations, while quantification involves estimating the number and capacity of panels, both of which demand high precision and reliability. To overcome these limitations, we investigate fine-tuning strategies within our PVAL framework that complement the prompting techniques. By fine-tuning the LLM on specialized datasets that include annotated satellite images with detailed spatial information, the model can better understand and perform multi-step reasoning required for localization and quantification.
@@
An additional benefit of PVAL is its capacity to provide nuanced likelihood and confidence values for identified solar panels. While conventional computer vision techniques have explored confidence scoring [24], LLMs can produce richer, context-aware assessments [25]. These confidence and likelihood metrics enable auto-labeling of large-scale datasets, reducing the need for manual annotation and enhancing system scalability. Consequently, powergrid operators can efficiently manage extensive data, ensuring accurate and reliable integration into ADNs and microgrids.
@@
In this work, we explore solar panel detection by proposing PVAL that combines large language models (LLMs) with data engineering, prompt engineering, and fine-tuning strategies. The key contributions of this paper are:
@@ 2 Problem Description
[…]
@@ 3 Methodology
Figure 2: Workflow for solar panel detection, localization, and counting using LLMs. The framework combines data engineering, prompt engineering, and fine-tuning to improve accuracy, with outputs including predictions with likelihood and confidence values.
@@ 3.1.1 Data Collection
The process begins by identifying geographic coordinates of solar panel installations from OpenStreetMap (OSM) via the Overpass API. These coordinates guide the retrieval of high-resolution satellite imagery using the Google Maps Static API. To ensure representativeness, we target diverse U.S. regions: Seattle, WA; Orlando, FL; Osage Beach, MO; Harlem, NY; Tempe, AZ; and Santa Ana, CA. Capturing a wide array of climatic conditions, urban densities, and architectural styles enhances the model’s ability to generalize across heterogeneous environments (see Figure 3).
@@ 3.1.2 Image Slicing
To enhance detection granularity and scalability, we divide each high-resolution satellite image into a $\times 4$ 4 grid, resulting in 16 equally sized tiles. The systematic slicing increases the number of training examples while preserving image resolution. Smaller solar panels remain distinguishable, enabling the model to identify, localize, and count panels at a finer spatial scale.
@@ 3.1.3 Data Labeling
We employ manual annotation to label each tiled image, recording the presence, location, and estimated quantity of solar panels. The human-in-the-loop approach ensures high-quality ground truth labels, facilitating effective fine-tuning and reliable performance evaluations [26].
@@ 3.3.3 Few-shot Prompting
Prompt – Few-shot #\## Example 1 (Solar) :
@@ 3.5 Likelihood and Confidence Mechanism for Auto-Labeling
Likelihood quantifies the probability that solar panels are present in the imagery. Higher likelihood values indicate stronger evidence supporting the presence of solar panels. Specifically, values approaching 111 suggest a strong match to features characteristic of solar panels, implying a high probability of their presence. Conversely, values closer to 0 denote a weak or negligible match, suggesting their absence.
Confidence reflects the language model’s internal certainty regarding its identification. Confidence values near 111 signify that the model is highly assured in its prediction, while values closer to 0 indicate greater uncertainty about the decision.
@@
4 Simulation and Analysis
@@
4.1 Experiment Preparation
@@
To rigorously assess the PVAL framework, we conducted experiments in two computational environments: The first environment utilized OpenAI’s platform for fine-tuning the GPT-4o-2024-08-06 model, allowing direct access to the service’s proprietary architecture and pre-trained capabilities. In parallel, we employed a high-performance computing server equipped with NVIDIA A100-SXM4-80GB GPUs to train and evaluate baseline models and perform additional analyses.
Our dataset comprises more than 100,000100,000100 , 000 satellite images of residential rooftops, each meticulously annotated to indicate the presence of solar panels. For a subset of the test samples, we also provided detailed spatial labels and estimated panel counts. All annotations were produced through a human-in-the-loop process, ensuring a reliable ground truth for both fine-tuning and validation. Figure 3 presents random samples from various regions.
4.2 Benckmark Models and Evaluation Metrics

### v2

@@ 1 Introduction
The global transition toward renewable energy has significantly influenced the operation and planning of power systems, particularly in active distribution networks (ADNs) and microgrids [1]. As a leading source of clean energy, solar photovoltaic (PV) technology has become increasingly important for decentralizing power generation and enhancing grid resilience [2]. To realize these benefits, accurate detection and mapping of solar PV installations in satellite imagery are critical for efficient energy management, demand-response optimization, and the seamless integration of distributed energy resources (DERs) [3]. However, the widespread adoption of solar panels has led to an exponential increase in the amount of imagery data, necessitating highly scalable and adaptable approaches for PV identification [4, 5].
@@
Early approaches to solar panel detection from imagery relied on traditional machine learning models, such as Logistic Regression [6], Random Forest [7], and Support Vector Machines based methods [8], often employing handcrafted features like texture and color. Although these methods were relatively efficient and required fewer training samples, they lacked the capacity to generalize to diverse spatial patterns. More recent research has turned towards deep learning [9, 10], particularly Convolutional Neural Network (CNN)-based models [11], which have demonstrated superior performance in extracting vision features from imagery data; recent advances span robust PV module segmentation and city-scale classification [12, 13, 14, 15]. Notable efforts such as Deep Solar [16] showcased large-scale solar panel mapping using Inception-v3, while U-Net [17] enabled global solar mapping. However, these methods often require substantial annotated datasets [18, 19]—a challenge amplified by roof superstructures and geographic heterogeneity [4]—and thus motivate label-efficient strategies [20, 21], intensive computational resources, and frequent retraining to adapt to changing conditions in different geographical regions and environmental contexts.
@@
These limitations highlight a pressing need for flexible, transparent, and resource-efficient solutions. Large language models (LLMs) have emerged as a promising paradigm for detection tasks with limited labels [22], owing to their extensive pre-training on diverse data sources. Although originally developed for natural language processing (NLP), advanced LLMs such as OpenAI’s GPT series [23], Google’s PaLM [24], and Meta’s LLaMA [25] have extended their capabilities to multimodal tasks, including image analysis. In remote sensing, vision–language models (VLMs) and datasets are rapidly maturing—e.g., RemoteCLIP[26], RS5M[27], and RSGPT [28]—alongside broad evaluations and surveys of multimodal LLMs [29, 30]. LLMs can capture rich contextual information [31, 32, 33] and handle heterogeneous inputs and varying contexts, making them particularly suitable for the dynamic environments typical of ADNs and microgrids [34].
@@
Although LLMs show great promise for solar panel detection, they face several challenges. LLMs often struggle with multi-step logical processes essential for accurately identifying and assessing solar panels [35], including top-view spatial reasoning required by overhead imagery [36]. These challenges stem from sequential disambiguation of confounders, spatial reasoning, and context integration in overhead scenes. Additionally, LLMs can exhibit inconsistencies in output formatting [37], which can lead to increased manual intervention and reduced scalability of solar energy management workflows.
@@
Another challenge is the misclassification of visually similar objects. For example, shadows, parking lots, or other flat surfaces may be incorrectly identified as solar panels due to their comparable appearance in satellite images. This issue is exacerbated in environments with high visual complexity or varying lighting conditions, which can obscure or distort the features LLMs rely on for accurate classification; related remote-sensing studies underscore the difficulty of robust shadow handling [38]. These failure modes motivate the task decomposition, schema-standardization, and few-shot strategies introduced below.
@@
To address the challenges associated with LLMs in solar panel detection, we propose prompting and fine-tuning strategies. We operationalize these strategies within an end-to-end PVAL pipeline. To manage this complexity, we implement task decomposition within our PVAL framework. We begin by defining the overall objective in the prompt, guiding the model to understand the context and goals of the analysis. The task is then broken down into specific steps: first, the model performs a holistic image analysis to detect potential solar panels; second, it pinpoints the exact locations of these panels within predefined regions; and third, it estimates the quantity of panels based on their arrangement and visibility. This structured approach ensures that each sub-task is handled systematically, reducing the likelihood of errors and improving the overall accuracy of detection.
@@
Second, consistent and interpretable outputs are crucial for integrating detection results into existing solar energy management systems (EMS). To achieve this, we employ output standardization by defining a JSON-based response format. This standardized structure includes key fields such as “solar panels present”, “location”, and “quantity”, each with predefined possible values to ensure uniformity. For example, the “location” field uses a detailed classification schema that divides the image into specific regions like “top-left” or “center” as illustrated in our visual schema. This standardization addresses the problem of inconsistent output formats, facilitating seamless data integration into geographic information systems (GIS) and energy management platforms. By ensuring that all outputs adhere to a uniform format, we eliminate the need for additional preprocessing, thereby streamlining workflows and enhancing the scalability of our detection system. In practice, this schema reduces format drift and makes outputs directly machine-actionable for downstream GIS/EMS pipelines.
@@
Third, accurate classification of solar panels, especially when distinguishing them from visually similar objects like shadows or parking lots, is a significant challenge. To address this issue, we utilize few-shot prompting by incorporating a limited number of highly relevant examples within the prompt. These examples include both positive cases (where solar panels are present) and negative cases (where they are absent). By using five-shot prompting, we enhance the model’s ability to generalize from these examples without extensive labeled datasets, consistent with findings in label-efficient remote sensing and multimodal few-shot learning [20, 21, 39]. We also curate exemplars that mirror common confounders (e.g., shadows, HVAC units) to improve robustness. This approach mitigates misclassification issues, enabling the LLM to better distinguish between similar-looking objects and thereby increasing the reliability of solar panel detection.
@@
Accurate spatial localization requires precise mapping of panel locations, while quantification involves estimating the number and capacity of panels, both of which demand high precision and reliability. To overcome these limitations, we investigate fine-tuning strategies within our PVAL framework that complement the prompting techniques. By fine-tuning the LLM on specialized datasets that include annotated satellite images with detailed spatial information, the model can better understand and perform multi-step reasoning required for localization and quantification. Fine-tuning complements prompting by aligning the model with rooftop-specific patterns and spatial cues.
@@
An additional benefit of PVAL is its capacity to provide nuanced likelihood and confidence values for identified solar panels. While conventional computer vision techniques have explored confidence scoring [40], LLMs can produce richer, context-aware assessments [41]. These confidence and likelihood metrics enable auto-labeling of large-scale datasets, reducing the need for manual annotation and enhancing system scalability; they also connect to contemporary work on calibration and uncertainty estimation in deep learning and NLP [42, 43, 44]. Operationally, we use these scores to gate auto-labeled samples and to prioritize uncertain tiles for targeted review. Consequently, powergrid operators can efficiently manage extensive data, ensuring accurate and reliable integration into ADNs and microgrids.
@@
In this work, we explore solar panel detection by proposing PVAL, which combines large language models (LLMs) with data engineering, prompt engineering, and fine-tuning strategies. The key contributions of this paper are:
@@ 2 Problem Description
Formally, let the input satellite image be represented as $I$. The objective is to develop a function $f(I)$ that maps the input image to a set of outputs $\{\textit{D},\textit{L},\textit{Q},\textit{C}\}$, where D indicates the binary detection of the presence of solar panels, L specifies the spatial localization of panels within the image $I$, Q represents the quantification of the number of solar panels in the image, and C provides a credibility score representing the likelihood of correct detection, where $C\in[0,1]$. The solution must ensure robust performance across diverse geographical regions, rooftop configurations, and varying environmental conditions. Additionally, the outputs $\{\textit{D},\textit{L},\textit{Q},\textit{C}\}$ must be interpretable and seamlessly integrable into decision-making frameworks for active distribution networks (ADNs) and microgrids.
@@ 3 Methodology
Figure 1: Workflow of PVAL: solar panel detection, localization, and counting using LLMs. The framework combines data engineering, prompt engineering, and fine-tuning to improve accuracy, with outputs including predictions with likelihood and confidence values.
@@ 3.1.1 Data Collection
We obtain geographic coordinates of solar panel installations from OpenStreetMap (OSM) using the Overpass API. [^1] https://overpass-api.de/ For each region, we preserve the full Overpass query, server endpoint, and export timestamp, which together define the exact OSM snapshot used. These coordinates serve as anchors for retrieving satellite imagery from the Google Maps Static API [^2] https://developers.google.com/maps/documentation/maps-static/overview, with parameters fixed at maptype=satellite, zoom=20, and size=400x400. Google imagery is publicly accessible with an API key, but redistribution is restricted by licensing. To ensure reproducibility, we release all metadata—coordinates, queries, request parameters, timestamps, and per-image hashes—so that the dataset can be regenerated by any user with their own API key. We target diverse U.S. regions (Seattle, WA; Orlando, FL; Osage Beach, MO; Harlem, NY; Tempe, AZ; and Santa Ana, CA) to capture a wide spectrum of climatic conditions, urban densities, and architectural styles, thereby enhancing the model’s generalization ability.
@@ 3.1.2 Image Slicing
Each high-resolution image is divided into a fixed $4\times 4$ grid, yielding 16 tiles per rooftop scene. This systematic slicing not only expands the training corpus but also preserves resolution at a finer spatial scale. Small-scale photovoltaic arrays remain visible, enabling the model to detect presence (D), localize placement (L), and estimate quantity (Q). The slicing process is deterministic, and we release both the slicing script and the metadata schema so that identical tile indices can be reconstructed.
@@ 3.1.3 Data Labeling
We employ manual annotation to label each tiled image, recording the presence, location, and estimated quantity of solar panels. Four annotators participated in the labeling. All annotators had prior experience in rooftop PV identification and visual interpretation of overhead imagery, and they received dedicated training before the annotation process. For every image, at least two annotators labeled independently, and we observed 100% pairwise agreement among all annotators, corresponding to a Kappa score of $1.0$. The human-in-the-loop approach ensures high-quality ground truth labels, facilitating effective fine-tuning and reliable performance evaluations [45].
@@ 3.3.3 Few-shot Prompting
Prompt – Few-shot #\# Example 1 (Solar) :
@@
A complete prompt Identify the presence of solar panels in images of residential rooftops, and determine their locations and quantity within the images. You will be provided with images that may contain residential rooftop solar systems. Analyze each image to detect solar panels.
Steps:
1. **Image Analysis**: Examine the entire image to identify any objects that appear to be solar panels.
2. **Panel Location**: Determine the coordinates or area within the image where the solar panels are located.
3. **Panel Quantification**: Calculate or estimate the number of solar panels based on their appearance and arrangement.
The output should be in JSON format, structured as follows, with each field restricted to specific possible values for consistency and accuracy:
"location": A description or coordinates indicating where the panels are located within the image. Possible values: [left, right, bottom, top, top-left, top-right, bottom-right, bottom-left, center, NA]
"likelihood_of_solar_panels_present": A quantitative measure of how probable it is that a certain hypothesis or model predicts the observed data. The value should be a decimal range from 0.0 to 1.0. 0.0 means a low likelihood, and 1 means a high likelihood.
Possible values for the likelihood of solar_panels_present are from 0 to 1.
"confidence_of_solar_panels_present": The strength of the model’s belief in its prediction. The value should be a decimal range from 0.0 to 1.0. 0.0 means a low confidence for the answer, and 1 means a high confidence for the answer.
Possible values for the confidence of solar_panels_present are from 0 to 1.
The output should have the following style:
# Example 1 (Solar):
{ "solar_panels_present": true,
"location": "top-left",
"quantity": "0 to 1",
"likelihood_of_solar_panels_present": 0.98,
"confidence_of_solar_panels_present": 0.90 }
# Example 2 (No Solar):
{ "solar_panels_present": false,
"location": "NA",
"quantity": "NA",
"likelihood_of_solar_panels_present": 0.21,
"confidence_of_solar_panels_present": 0.87 }
@@ 3.5 Likelihood and Confidence Mechanism for Auto-Labeling
Likelihood quantifies the probability that solar panels are present in the imagery. Higher likelihood values indicate stronger evidence supporting the presence of solar panels. Specifically, values approaching $1$ suggest a strong match to features characteristic of solar panels, implying a high probability of their presence. Conversely, values closer to 0 denote a weak or negligible match, suggesting their absence.
Confidence reflects the language model’s internal certainty regarding its identification. Confidence values near $1$ signify that the model is highly assured in its prediction, while values closer to 0 indicate greater uncertainty about the decision.
@@
To summarize the overall workflow, we provide a step-by-step description of the pipeline with fine-tuning in Algorithm 1.
Input: Regions $\mathcal{R}$, Overpass query $\mathcal{Q}$, MapsKey, grid $G{=}4{\times}4$, schema $\mathcal{S}$, few-shot $\mathcal{E}$, temperature T=0T{=}0; thresholds $(\tau_{p},\tau_{c})$, OpenAI api_key; base_model (e.g., gpt-4o); hyperparams (epochs = 5)
Output: Fine-tuned model id $\theta^{\ast}$
[…]
2) Tile & seed labels: Slice each image into $G$ tiles; attach human labels $(D,L,Q)$ where available, else $\varnothing$.
3) Prompted inference (unlabeled tiles): For each unlabeled tile $t$: build $P=\text{AssemblePrompt}(t,\mathcal{E},\mathcal{S})$; run $\hat{y}=\text{GPT4o}(P,T)$.
[…]
5) Build fine-tuning set: $\mathcal{D}_{\text{ft}}=\{(t,y):y\neq\varnothing\}$; serialize JSONL lines (base64 image, prompt with $\mathcal{E}$ and $\mathcal{S}$, target JSON $\{D,L,Q\}$) as solar.jsonl.
6) Fine-tuning via OpenAI API:
(a) Initialize client): $\text{client}\leftarrow\text{OpenAI(api\_key)}$.
(b) Upload training file: $\text{file\_id}\leftarrow\text{client.files.create(file=solar.jsonl, purpose="fine-tune").id}$.
(c) Create job: $\text{job}\leftarrow\text{client.fine\_tuning.jobs.create(file\_id, base\_model, n\_epochs, batch, lr})$.
(d) Poll status): while status $\notin$ {succeeded, failed} do
  $\text{job}\leftarrow\text{client.fine\_tuning.jobs.retrieve(job.id)}$; status $\leftarrow$ job.status;
(e) On success): $\theta^{\ast}\leftarrow\text{job.fine\_tuned\_model}$; result_files $\leftarrow$ job.result_files.
7) Evaluate & use fine-tuned model: Evaluate on held-out set at T=0T{=}0 (det: P/R/F1; loc/qty: exact-match). For future predictions, reuse Steps 3–4 with model set to $\theta^{\ast}$.
return $\theta^{\ast}$.
Algorithm 1 PVAL Pipeline with Fine-Tuning Process
4 Experiments setup and Preparation
@@
4.1 Experiment Setup
To rigorously assess the PVAL framework, we conducted experiments in two complementary computational environments. The first environment utilized OpenAI’s platform for fine-tuning the GPT-4o-2024-08-06 model, which provided direct access to the proprietary multimodal architecture and its pre-trained capabilities through the official API. In parallel, we employed the Sol supercomputer at Arizona State University [53], a high-performance computing cluster designed for large-scale AI research. The Sol environment featured NVIDIA A100 GPUs with 80 GB of memory (CUDA 12.7) and dual AMD EPYC 7413 processors with 48 CPU cores running at 3.6 GHz, supported by a Linux operating system (kernel 4.18). Each job was executed on compute nodes with sufficient GPU memory to process large batches of rooftop imagery efficiently. All baseline models and additional simulations were implemented in Python (version 3.9) using the PyTorch framework, with development and debugging conducted in the PyCharm IDE. This setup ensured both high computational throughput and reproducibility across experiments.
From a computational efficiency perspective, baseline CNN and ViT models required significant training compute on the Sol cluster (approximately 15–40 GPU-hours depending on architecture), whereas prompted PVAL required no local training and fine-tuned PVAL incurred only a one-time cloud-managed training cost. Inference with PVAL takes longer per image (about 1–3 seconds via API) compared to CNN/ViT baselines on A100 GPUs (0.2–0.5 seconds). However, PVAL eliminates repeated retraining when expanding to new regions or schema modifications. This trade-off demonstrates reduced overall resource requirements by shifting the cost from local GPU training cycles to a lightweight API-based workflow.
@@ 4.1 Experiment Setup
4.2 Data Materials
Data sources. Our dataset combines U.S. and global satellite imagery to comprehensively evaluate solar PV detection. For U.S. regions, we collected data from six representative locations: Seattle (WA), Orlando (FL), Osage Beach (MO), Harlem (NY), Tempe (AZ), and Santa Ana (CA). To assess cross-continental generalization, we further extended the study to six globally distributed cities spanning diverse climate zones and urban morphologies: Sydney (Australia), Cape Town (South Africa), Kuwait City (Middle East), Oxford (Europe), São Paulo (South America), and Shanghai (Asia). Geographic coordinates of PV installations were obtained from OpenStreetMap (OSM) via the Overpass API, and high-resolution imagery was retrieved from the Google Maps Static API.
Transparency of data. Existing methods often lack transparency regarding their underlying algorithms or training datasets, as their annotated PV datasets are not publicly available. By contrast, our dataset is constructed entirely from openly accessible sources—geographic coordinates from OSM and imagery from Google Maps—making the collection process transparent and reproducible. Although annotation requires manual effort, the underlying imagery is straightforward to obtain, ensuring that our dataset can be independently replicated and extended by other researchers.
Data structure. Each rooftop image was divided into a $4\times 4$ grid, producing 16 tiles per rooftop. This tiling strategy expanded the dataset size while preserving spatial resolution, ensuring that small-scale PV installations remained visible. All tiles were labeled with the binary field solar panels present. A subset of $10,000$ tiles was further annotated with detailed labels for location and quantity. The location label follows a predefined schema including top, bottom, left, right, center, and quadrant-based categories (e.g., top-left, bottom-right). The quantity label records the estimated range of panel counts using discrete intervals: $(0,1]$, $(1,5]$, $(5,10]$, and $10,+\infty)$. As illustrated in Figure 4, these structured annotations provide a consistent representation of rooftop PV characteristics.
Figure 4: Illustration of the dataset structure. Each rooftop image is divided into tiles and labeled with solar panel presence, location and quantity information. Example entries are shown with corresponding annotations.
Dataset summary. Table 1 provides an overview of the datasets. For consistency, both baseline models (VGG, ResNet, U-Net, ViT, etc.) and our proposed PVAL framework were trained on the same fine-tuning dataset, which consists of 2,000 annotated rooftop tiles from Santa Ana, CA. The large-scale evaluation set contains approximately 100,000 tiles from Tempe and Santa Ana. Additional datasets from the remaining U.S. and international regions, each with 480 tiles, were used to assess cross-regional and cross-continental generalization.
Table 1: Summary of training and evaluation datasets. All baseline models use the same fine-tuning data as PVAL.
Region / City Role Size
Santa Ana, CA Training (fine-tuning) 2,000 tiles
Tempe, AZ + Santa Ana, CA Large-scale Test $\sim$100,000 tiles
Seattle, WA Cross-regional Test 480 tiles
Orlando, FL Cross-regional Test 480 tiles
Osage Beach, MO Cross-regional Test 480 tiles
Harlem, NY Cross-regional Test 480 tiles
Sydney, Australia Cross-continental Test 480 tiles
Cape Town, South Africa Cross-continental Test 480 tiles
Kuwait City, Middle East Cross-continental Test 480 tiles
Oxford, Europe Cross-continental Test 480 tiles
São Paulo, South America Cross-continental Test 480 tiles
Shanghai, Asia Cross-continental Test 480 tiles
4.3 Benchmark Models